OpenStudy (anonymous):

If the linear equation 4x = 10 − 5y is written in standard form, what is the value of B? Numerical Answers Expected!

OpenStudy (anonymous):

4x = 10 - 5y 5y =10 - 4x y = -4/5 x + 2 b is 2.